What would be the linear algebraic equation with clearly defined variables to represent the cost of each option?
My scenario is comparing the cost options of a yearly pass to a theme park which is $65.00, to daily visits, which cost $15.00 daily. The park will be visited 12 times during the year.

Answer by richwmiller(17219)

y=65
y=15x where x=12 visits
y equals annual cost
y=15*12=5*3*3*4=9*20=180
y=15x-65 would be the difference in costs with x the number of visits and y the difference in the two options